Loving Gardens has $6 million in assets, $700,000 EBIT, 80,000 shares of stock outstanding, and a marginal tax rate equal to 40 percent. If Loving Garden's debt-to-total-assets (D/TA) ratio is 70 percent, it pays 12 percent interest on debt, whereas if the D/TA ratio is 40 percent, interest is 9 percent. Calculate Loving Gardens EPS and ROE (ROE = Net Income / Equity) for each capital structure. Which capital structure is better?
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

E F G Total Assets D/TA ratio Debt Equity Interest rate on debt $6,000,000.00 70% $4,200,000.00 $1,800,000.00 12% Total Asset

Cell reference -

A B C D Total Assets D/TA ratio Debt Equity Interest rate on debt 6000000 0.7 =C2*C3 =C2-C4 0.12 Total Assets D/TA ratio Debt

Please note: It assumed that share outstanding is same for both cases.

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.